OpenStudy (anonymous):

9 5/6 improper fraction

OpenStudy (anonymous):

just multiply the denominator with the nine and then add the numerator so 6 x 9 = 54 + 5 = 59/6

OpenStudy (anonymous):

The answer is 59/6

OpenStudy (anonymous):

9x6+5 54+5 59/6